OpenStudy (anonymous):

@satellite73 @bibby Solve for x. Enter your answer below as a fraction in lowest terms, using the slash ( / ) as the fraction bar. x+(5/12)=5/8

OpenStudy (anonymous):

subtract \(\frac{5}{12}\) from both sides, i.e. compute \[\large x=\frac{5}{8}-\frac{5}{12}\] now it is basically an arithmetic problem

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

least common multiple of 8 and 12 is 24, so compute \[x=\frac{15}{24}-\frac{10}{24}\]